ON RATES OF CONVERGENCE OF INFORMATION THEORETIC CRITERION IN RANK DETERMINATION OF ONE-WAY RANDOM EFFECTS MODELS

Y. WU

Department of Mathematics and Statistics, York University, 4700 Keele Street, North York, Ontario, Canada M3J 1P3

(Received February 28, 1992; revised September 2, 1992)

Abstract. In this paper, upper bounds on the probabilities of wrong determination of the rank of covariance matrix of random effects in one-way random effects models are given, based on the information theoretic criterion. Under weak conditions, the bounds are shown of exponential-type.

Key words and phrases: Rate of convergence, information theoretic criterion, one-way random effects model, rank determination.